Abdul recently got a new job selling artisanal pickles at farmers' markets. He's paid $20 per hour, and works 40 hours a week. He quite enjoys the work, but driving in traffic is a pain, so his average disutility from work is $2 per hour.


a. Now, suppose that Abdul expects it will take him 6 weeks to find a new job if he loses this one. Calculate his total employment rent. Describe in words what Abdul's employment rent represents. What does it tell us about the amount of effort he is likely to put in?


b. Now, suppose Abdul gets a new manager who is very difficult to work with, increasing his average disutility from work to $5 per hour. What is his new total employment rent? How would you expect his behavior to change?


1
Expert's answer
2021-03-29T11:34:16-0400

a) "20-2=18" hourly employment rent

"(20-2)*40*6=4,320" - It represents the amount that he will lose while he is out of work. Which is reduced by the amount of work-related disutility. He spent ten percent of his work-related efforts on disutility.

b) "20-5=15" new hourly employment rent. He will be more eager to find a new job.